### v1

# Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals

Abstract
Here we present the discovery and characterization of TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 family. We previously reported on the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) kagome metals. This structure type represents a magnetically diverse and exfoliable family of kagome metals. Detailed characterization of the magnetic and electronic transport properties of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state. The system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (-50%) near a metamagnetic transition. Unlike the Lu-Gd analogs, TbTi3Bi4 cannot be synthesized by the methodology presented previously due to phase competition with another, unreported class of materials, the Ln2-xTi6+xBi9 (x1.7–1.2) (Ln: Tb–Lu) shurikagome metals. Like the LnTi3Bi4 kagome metals, the Ln2-xTi6+xBi9 family exhibits quasi-2D networks of titanium and chains of rare-earth. We present the structures and some select basic properties of the Ln2-xTi6+xBi9 family alongside the newly discovered TbTi3Bi4. A discussion regarding the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families helps frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and provides a strategy to circumvent formation of Ln2-xTi6+xBi9. 111Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
I Introduction
Spurred on by continued discoveries in the AM3X5, AM6X6, and CoSn-type compounds, kagome metals have proven to be a rich space for exploring the intersection of solid-state chemistry and fundamental condensed matter physics. The impetus for much of the inspiration comes from the basic electronic structure afforded by the kagome motif – one which can contain Dirac points, flat bands, and Van Hove singularities. [1, 2, 3, 4]. Derived from the triangular lattice, the kagome motif have shown remarkable structural and chemical diversity, allowing for a rich interaction between structural chemistry, electronic properties, and magnetism. As a prime example, the AM3X5 compounds have been heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, superconductivity).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity.[14] In fact, embedding magnetism into the kagome motif underpins much of the research into the AM6X6 and CoSn-type compounds, which continues to provide the motivation for the development of new kagome materials.[15, 16, 17, 18, 16, 19, 20, 21, 22, 23, 24, 25, 26]
One fundamental way to develop new materials is to perturb simpler, “fundamental” prototype structures. Take, for example, the CoSn family of compounds. By adding a filler atom in the natural voids formed by the structure, the CoSn-type stacking gives way to another of the most researched kagome families, the AM6X6. Further perturbations on the AM6X6 prototype yield a grandchild of the CoSn prototype – the AM3X4 kagome metals.
We reported previously on the discovery of both the antimonide AV3Sb4 (A: Eu2+, Yb2+) [27] and bismuthide ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[28] which make up a larger family of AM3X4 compounds.[27, 28, 29, 30, 31, 32, 33] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals. The A-site elements comprise zig-zag chains which run parallel to the M-site kagome layers. Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds maintain a much more quasi-2D crystal structure. Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure. Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission (ARPES).[32, 34, 35, 36, 37, 38]
Under the conditions presented in our previous work, LnTi3Bi4 compounds containing rare-earth atoms smaller than Gd3+ did not appear. In this work, we present a modified growth procedure which allows stabilization of TbTi3Bi4, the smallest rare-earth cation reported in the LnTi3Bi4 structure thus far. Our investigation into the smaller rare-earth species (Tb–Lu) also unveiled an intimate competition between the LnTi3Bi4 kagome metals and a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals. By understanding the phase diagram surrounding the LnTi3Bi4 compound, we are able to subvert the formation of Tb2-xTi6+xBi9 and stabilize single crystals of TbTi3Bi4. This work details our rationale for the stabilization of TbTi3Bi4 alongside a discussion of its magnetic and electronic transport properties. Remarkably, TbTi3Bi4 presents some of the most rich and complex magnetism in the known AM3X4 compounds, reinforcing the need to understand and control the formation of new phases. Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals, which have strong chemical disorder and unusual magnetic behavior.
II Methods
II.1 Single Crystal Synthesis
TbTi3Bi4 single crystals are grown through a bismuth self-flux. Tb metal (Ames), Ti (Alfa 99.9% powder), and Bi (Alfa 99.999% low-oxide shot) were measured in a 2:3:20 ratio into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit.[39] The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules. The flux was heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h. The sample was then quenched by removing it from the furnace and immediately placing in a preheated 600°C furnace. The growth proceeded with a cooling rate of 2°C/hr to 400°C, where excess bismuth was removed through centrifugation. Ln2-xTi6+xBi9 single crystals are grown in a similar manner, except the samples are cooled linearly from 1100°C to 600°C at a rate of 2°C/hr before centrifugation at 600°C.
II.2 Bulk Characterization
Single crystals were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD). Diffraction data were collected at 100 K on a Bruker D8 Advance Quest diffractometer with a graphite monochromator using Mo K radiation ( = 0.71073 Å). Data integration, reduction, and structure solution was performed using the Bruker APEX4 software package. As noted previously, diffraction on LnTi3Bi4 crystals is challenging and requires face-indexing absorption corrections.[28] All atoms were refined with anisotropic thermal parameters. CIF files for all structures are included in the ESI. Laue diffraction utilizing a Photonic Science vertical Laue system was used to orient crystals of TbTi3Bi4 for physical property measurements.
Magnetization measurements of LnTi3Bi4 single crystals were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode. Additional measurements below 1.8 K on the MPMS3 utilized an additional He3 insert. Angle-resolved magnetization measurements were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
Heat capacity measurements were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS) equipped a dilution refrigerator insert. Thermal contact was achieved using Apezion N-grease. Electronic resistivity measurements were performed in a Quantum Design 9 T Dynacool PPMS. Single crystals were mounted to a sheet of sapphire and then the sapphire plate was then adhered to the sample puck stage using GE varish. The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane. Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%).
III Results and Discussion
III.1 Crystal Structure and Synthesis
Our prior work identified that the ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[28] However, growth of ATi3Bi4 crystals containing smaller rare-earth metals (Tb–Lu) was stymied due to preferential formation of Ln2-xTi6+xBi9. Figure 1(a) provides a simple schematic showing the lanthanide row sorted in order of the 9-coordinate Shannon ionic radius. Within the ATi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state. As such, the radii of Eu2+ and Yb2+ are analogous to the larger La3+ and Pr3+, respectively. Large cations preferentially stabilize the ATi3Bi4 phase (denoted by blue), and small cations preferentially stabilize Ln2-xTi6+xBi9 (green). Within this work, we demonstrate that Tb represents a unique cross-over between the two phases, and both TbTi3Bi4 and Tb2-xTi6+xBi9 can be obtained by tuning the growth profiles.
Figure 1(b,c) details the crystal structure of both the ATi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds. As we have discussed ATi3Bi4 extensively before, for simplicity we opt to show only half the unit cell here (the omitted half is generated through an inversion operation, space group Fmmm. ATi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln-Ln zig-zag chains. The local coordination of the Ln species is approximately 9-fold. Crystals of ATi3Bi4 are quasi-2D and can be easily exfoliated using adhesive tape. We have observed that both M-Sb and Ln-Sb terminations can be achieved.[27] The kagome network in ATi3Bi4 compounds is only slightly distorted, though the quasi-1D zig-zag chain limits the symmetry of the system to orthorhombic.
The Ln2-xTi6+xBi9 compounds can be understood as a heavily alloyed form of Ti8Bi9[40], where Ln atoms preferentially occupy two of the Ti sites. If the substitution was perfect, (e.g. Ln2Ti6Bi9) the Ln atoms would create a nearly ideal 1D chain that cuts perpendicular to the Ti–Ti network. The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to ATi3Bi4 (4Å)The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles. Paying homage to both the linguistic origin and coordination of the “kagome lattice,” the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
The two sites that are preferentially occupied by Ln atoms possess considerably different coordination than the remaining Ti sites in the parent Ti8Bi9 structure. Considering bonds 3.2Å, Ti atoms within the shurikagome network have an approximate coordination of 10 (4 Ti–Ti bonds, 6 Ti–Bi bonds) and a polyhedral volume of approximately 50.7Å3. Sites preferentially occupied by Ln possess a coordination of 8, solely from Ln–Bi bonds, and form slightly distorted square anti-prisms. The two mixed Ln/Ti sites are also distinct from one another, and one has a substantially larger polyhedral volume (56.9Å3) than the other (49.5Å3). As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and further preferentially occupy the largest of the mixed Ln/Ti sites. Generally speaking, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site. As such, the total stoichiometry can be more precisely written as Ln2-xTi6+xBi9 (x1.7–1.2). CIF files for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI. We will discuss trends in the Ln substitution alongside the magnetic properties in a later section.
As alluded to before, TbTi3Bi4 is not stable under conditions presented in our first manuscript.[28] Samples of large rare-earth ATi3Bi4 compounds are readily grown through a simple flux (cooled from 1100°C to 600°C at a rate of 2°C/hr). However, samples containing Tb will exclusively form Tb2-xTi6+xBi9. Figure 2(a) shows a small section of the Tb–Ti–Bi phase diagram near several of the phases of interest. Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[41, 42] Note the proximity of TbTi3Bi4 and Tb2-xTi6+xBi9 (particularly the proximity of the the hypothetically ideal Tb2Ti6Bi9). The large single-phase region of Ti8Bi9 (and the associated alloys with Tb2-xTi6+xBi9) is problematic, as it isolates a large swath of the phase diagram from Bi-rich fluxes. The single phase region even persists beyond where a hypothetical Bi2Ti6Bi9 exists (complete substitution of Bi on the mixed sites). There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[43] However, if Tb2Ti6Bi9 is not stable, then TbTi3Sb4 should still be accessible from a Bi self-flux.
We designed a series of experiments to test both the thermal stability of Tb2-xTi6+xBi9 and compositional dependence. We first observed that samples allowed to completely freeze (T=270°C) often exhibited Tb2-xTi6+xBi9 crystals that were rounded, as though they started to decompose. Modifying the flux towards Tb-rich conditions instead resulted in rounded crystals of TbBi. In both cases, the corners of the decomposing crystals were encased in agglomerates of TbTi3Bi4. This highlighted that TbTi3Bi4 is a low-temperature phase. Subsequent quenching experiments further suggested that TbTi3Bi4 is likely only stable below 600°C. Samples grown above 600°C reveal two alkemade triangles; 1) three phase TbBi–Tb2-xTi6+xBi9–Bi and 2) two phase Tb2-xTi6+xBi9–Bi. Samples grown below 600°C reveal instead; 1) three phase TbBi–TbTi3Bi4–Bi and 2) three phase TbTi3Bi4–Tb2-xTi6+xBi9–Bi. Figure 2(b) presents a simple projection of Figure 2(a) along the TiBi–TbBi pseudobinary line, summarizing the relative proximity and temperature stability of these results.
Crystals of TbTi3Sb4 are 1-5 mm pseudo-hexagonal crystals with a brilliant metallic luster. They are readily exfoliated and are lightly air-sensitive. Samples can be worked with in air, though they will swell and spall under extended (1 day) exposure. Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster. Samples are readily cleaved with a blade parallel to the ab-plane. Ln2-xTi6+xBi9 crystals are substantially more sensitive to air, and crystals will completely disintegrate within a few hours. In both cases, the samples expand and self-cleave along the ab-plane upon extended exposure to air.
III.2 Properties of TbTi3Bi4
Due to the reduced symmetry (Fmmm), the ATi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains. TbTi3Bi4 is no exception. To diagnose the potential anisotropies in the material, we first begin with a study of the magnetization along different principle crystallograpic axes. Figure 3(a) presents the magnetization of a single crystal of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two primary transitions at 24 K and 3 K. The [100] direction corresponds to the “easy axis,” and orienting the magnetic field parallel to the [010] or [001] induces a substantially weaker effect. The behavior is reminiscent of GdTi3Bi4[28], though crystals of TbTi3Bi4 respond negligibly to fields oriented along [001].
Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions. Upon cursory inspection, the data again resembled that of GdTi3Bi4 (two metamagnetic transitions when ). However, the metamagnetic transitions in TbTi3Bi4 exhibit substantial field hysteresis. The data saturates at approximately The first transition plateaus at approximately 9.2 , in good agreement with the 9 expected for Tb3+. The first metamagnetic transition first plateaus at a moment of 2.9-3.2 , which is approximately 1/3 of the expected net moment. Furthermore, closer inspection at the second metamagnetic transition reveals an unusual inflection midway through the rising field direction (see Figure 3(b,inset)). This inflection occurs at approximately 6.2 slightly above 2/3 the expected moment. This inflection and other properties of the isothermal magnetization curves will be studied in detail below. Similar to the temperature-dependent magnetization, there are only weak, linear responses when or . Figure 3(c) provides a series of magnetization curves collected while rotating between , , and in 15°increments. The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further easy-[100]) anisotropy.
Figure 3(c) presents the inverse susceptibility, collected under a 100 Oe field where . The high temperature data exhibits a slight upwards curvature, potentially due to a temperature-independent contribution from core diamagnetism (bismuth) or the quartz sample holder. We show Curie-Weiss fits to the data with (gray) and without (black) a temperature-independent term. While the gray curve fits over a longer range of data, it requires a large emu mol-1 Oe-1 and yields an unphysical effective moment of 11 . We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit, which yields a =53 K and a = 9.9. The effective moment agrees well with that expected of Tb3+ (9.72), though the Weiss temperature is unphysically large (53 K), likely affected by the temperature independent contribution, limited temperature range, and crystal-field effects.
As a consistency check with the magnetization data, we performed zero-field heat capacity measurements down to 100 mK. The sample was initially oriented with such that measurements could also be made for the intermediate plateaus seen in Figure 2(b). The first order hysteretic nature of the transitions seen in the isothermal magnetization made the quantative analysis of field-dependent heat capacity challenging, and we have opted to reserve this data for a later study. Instead, we focus on the analysis of the zero-field heat capacity utilizing samples of LaTi3Bi4 as a nonmagnetic standard. Figure 3(d) demonstrates the experimental heat capacity data (green) as a function of temperature, plotted in . The two magnetic transitions are evident around 20 K and 2 K. At temperatures below 2 K, a nuclear Schottky peak dominates the heat capacity signal. The peak is still increasing at 0.2 K, which agrees relatively well with a Tb-based Schottky peak ( K).[44] Utilizing LaTi3Bi4 as the nonmagnetic standard (black) and invoking a simple nuclear Schottky model , we can extract the magnetic entropy for TbTi3Bi4.
Figure 3(e) shows the integrated magnetic entropy (green) as a function of temperature. The extracted magnetic contribution to the heat capacity () is also shown (gray) as a schematic reference for the magnetic transition temperatures (not to scale). Interestingly, integrating up to the primary magnetic transition appears to correspond with an entropy release of approximately . Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other ATi3Bi4 materials.[28] By 100 K we recover approximately 90% of the total entropy expected for Tb3+. Considering that the higher temerature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
We now return to the hysteretic metamagnetic transitions observed in Figure 3(b). Closer inspection into the second metamagnetic transition reveals a complex magnetic landscape which can only be elucidated with high density phase maps. Figure 4(a,d) present our efforts to map the magnetic landscape of TbTi3Bi4, demonstrating a set of isothermal magnetization curves collected in both the increasing (a) and decreasing (d) field directions. One can immediately see that the sharpness of the second transition belies a network of subtle changes in the magnetization. Taking the derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e). Immediately one can see that the plateau visible in the 1.8 K magnetization data shown before in Figure 3(b) fragments into a multitude of unique states at intermediate temperatures from 2–15 K. To better highlight the intricacy of the magnetic landscape near the potential 2/3 magnetization plateau, we collected higher density data over the limited ranges highlighted by the black dashed lines.
Figure 4(c,f) demonstrate the phase diagram for TbTi3Bi4 in the vicinity of the second metamagnetic transition. At least 5 unique phase pockets can be observed. Furthermore, while the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field. Data was also collected for the zero-field-cooled virgin curves, though they are qualitatively similar to Figure 4(a,b,c) and have been omitted for brevity. A series of line cuts (Figure 4(c,right)) through the phase diagram data (white dashed lines) helps to clarify the clear boundaries between the phase pockets. Considering the complexity of the magnetic landscape, neutron diffraction experiments will be invaluable to identify the types and natures of the individual magnetic phases.
Antiferromagnetic metals are not particularly common, particularly so for materials that can be readily exfoliated. While the low temperature magnetism of the rare-earth metals prevents facile applications within fields like spintronics, understanding how the magnetism interacts with the transport properties is key to understanding the connection between the kagome network (which dominates the electronic states near E) and the magnetism. Figure 5(a) shows a basic resistivity curve as a function of temperature under zero-field conditions. Our experiment passes current parallel to the [010] direction, and orients the field such that (oriented to induce the metamagnetic transitions). The crystal quality is good (RRR40) and the zero-field transport is indicative of a relatively high mobility metal. The magnetic transition at 20 K is evident in the resistivity, highlighted by a sharp cusp and subsequent drop. This is the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
We now apply a magnetic field to examine the magnetoresistance (MR) of TbTi3Bi4 in the same experimental setup. Figure 5(b) shows the symmetrized transverse magnetoresistance with and at 1.8 K. As one would expect from the magnetization results, we observe two sharp transitions in the magnetoresistance, one for each of the metamagnetic features. The magnetoresistance retains the qualitative field hysteresis of the magnetization data, both in magnitude and location. The first metamagnetic transition marks a sharp increase in the magnetoresistance of approximately 30–40%. The second metamagnetic transition marks a sharp decrease in the magnetoresistance to approximately 50%, which one may expect for a total polarization of the Tb moments. We find the sharp switching between positive and negative MR intriguing in TbTi3Bi4, particularly for a system which is already quite metallic (1Ohm-cm at 1.8 K).
Given the strong MR response in TbTi3Bi4, we were curious to investigate whether the complex landscape of magnetic transitions seen in the phase diagrams would influence the electronic transport. Figure 5(c) provides a series of magnetoresistance curves as a function of temperature from 2 K to 21 K in 1 K steps. For simplicity and clarity, we are only showing the data under increasing fields. We would like to highlight the behavior near the second transition of the symmetrized data (3 T). Immediately before the transition to negative magnetoresistance, we observe that the magnetoresistance cusp becomes a peak – then splits into two peaks – merge back together – and then vanish as the system approaches the Neel temperature. These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4(c). This further suggests that these are bulk effects corresponding to distinct types of magnetic order.
The sample geometry prevents us from measuring the Hall response where and . The in-plane moments would require thinning the sample perpendicular to the natural layers of the crystal, which typically causes the sample to flake and fall apart. However, we can readily orient the sample such that and to explore the Hall effect and magnetoresistance when the metamagnetic transitions are not induced. Figure 5(d) shows the antisymmetrized Hall resistivity for TbTi3Bi4 over two temperature regimes. The top figure shows the Hall response from 2 K to 18 K, and the bottom figure highlights temperatures above the magnetic transition temperature (20 K to 30 K). No subtractions or corrections have been performed besides antisymmetrization. All data sets exhibit a strongly nonlinear Hall response over all fields. A signal that looks deceptively like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature. Furthermore, at high temperatures we recover an additional inflection in the data. The signal is also relatively small, leading us to suspect that we may be seeing the effect of multiband transport. We can look to some prior work on NdTi3Bi4 for some guidance – since the electronic structure of these materials doesn’t appreciably change with the choice of rare-earth. Experimental ARPES data has observed a Dirac-like feature near the Fermi level[38, 37, 35, 36] and has further observed gapping of the Dirac-like bands as NdTi3Bi4 transitions into the magnetic ground state[35]. The sample exhibits strong positive magnetoresistance (200%) when and , in stark contrast to behavior when and .
Overall, our work into TbTi3Bi4 has revealed a unique antiferromagnetic, cleavable metal with an exceedingly complex landscape of magnetic phases. The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative magnetoresistance near the 2/3 magnetization plateau, and an unusual coupling between the magnetic phase pockets and the transport. We now turn to briefly investigate the magnetic properties of the Ln2-xTi6+xBi9 shurikagome metals and the influence of disorder on the Ln–Ln chains.
Shurikagome Ln | Occ | Occ | Tot Ln | Tot Ln | M () | M () | M () | T (K) | () | (K) |
Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.7 | -24 |
Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | -14 |
Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | -0.2 |
Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.6 | -0.9 |
Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 7.3 | 14 |
Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | 0 | 0 | 0.1 | N/A | N/A | N/A |
III.3 Properties of Ln2-xTi6+xBi9
Diffraction studies on the Ln2-xTi6+xBi9 compounds revealed several interesting observations. We first attempted to solve the structure using fully occupied Ln sites (forming fully occupied chains). While R values were actually quite favorable, there existed large amounts of excess electron density around the Ln04 and Ln05 sites. The residual Q-peaks can be eliminated by allowing the occupancy of the Ln sites to vary, which results in partially occupied Ln chains. However, compositions extracted from single-crystal X-ray diffraction (SCXRD) using the assumption of pure (although partially occupied) Ln sites are in strong disagreement with electron dispersive spectroscopy (EDS) results. Bluntly, there is not enough Ti in the unit cell derived from SCXRD. This can be remedied by allowing antisite disorder on the Ln sites, allowing the composition of the Ln to vary, but restraining it such that the site remains fully occupied (e.g. Ln04 + Ti04 = 1). Even this appears to slightly underestimate the amount of Ti, and EDS results suggest that some Ti may be substituting elsewhere as well. This is perhaps expected, as the published Ti–Bi phase diagram indicates a full solid solution over the range of 50:50 Ti:Bi to 35:65 Ti:Bi.[41]
However, the SCXRD data did not independently suggest any additional site substitutions, and the constraints mentioned above yielded quite favorable statistics. Figure 6(a) highlights the cell volume for the Ln2-xTi6+xBi9 shurikagome metals as a function of the 8-coordinate Shannon radius. Despite the variable rare-earth occupancy, the compounds still adhere to a nearly linear expansion in cell volume, as one would expect of a rare-earth series. These results are briefly summarized in Table 1. While CIF files are provided in the ESI, here we focus on the influence of the Ln atom and the disorder in crystals of Ln2-xTi6+xBi9. The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln04 and Ln05. As mentioned before, the Ti occupancy on those sites is constrained such that they add to 1. The sum of the SCXRD-determined occupancies is provided in the column labeled “Tot Ln.” This can be directly compared with the composition extracted from EDS measurements in the next column.
SCXRD appears to consistently calculate higher concentrations of Ln as compared to EDS measurements, over approximating by about 15%. Some interesting trends appear in the overall data. Both the Ln04 and Ln05 site occupancies decrease as we move to smaller Ln atoms. This effect is substantially more pronounced on the Ln05 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9. We performed some exploratory syntheses around Tb2-xTi6+xBi9 in an attempt to fully saturate the Ln site, though only small changes in the rare-earth composition were noted.
Considering the high disorder in crystals of Ln2-xTi6+xBi9, particularly within the Ln sublattices, it is not clear what magnetic response to expect. Naively, the most likely scenario was freezing into a spin-glass, or a complete loss of a well-defined magnetic ground state in the non-Kramers ions. We now turn to a brief overview of the magnetic properties of the Ln2-xTi6+xBi9 compounds.
The compounds are tetragonal (P4/nmm) yielding two directions of interest, (parallel to chain) and (perpendicular to chain). We did not consider rotations with in the ab plane for this work. All compounds exhibit weak anisotropy with the exception of Tm. When , Tm2-xTi6+xBi9 exhibits an order of magnitude stronger response. For Tb, Dy, and Ho, data collected with exhibits a cusp in the magnetization, strangely resembling an antiferromagnetic transition. For Er and Tm, the data with exhibits a broad peak instead. All other combinations of elements and directions (e.g. Tb with are featureless, although the magnitude of the magnetization remains comparable to the other direction). These results are summarized in Figure 6(a) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction. Please note that the magnetization results are shown in arbitrary units for ease of comparison. Full data sets are provided in the ESI.
Note that the data presented in Figure 6(a) includes both zero-field-cooled (ZFC) and field-cooled (FC) data. The lack of ZFC/FC splitting is surprising, as we may reasonably expect the strong disorder to result in a glassy magnetic state. Furthermore, AC susceptibility measurements on Tb2-xTi6+xBi9 indicate no frequency dependence. However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits a strong frequency response. It is striking that the features follow the temperature scale expected of the latter half of the rare-earth row – Tb possesses the highest ordering temperature, and Tm possesses the lowest.
Figure 6(b) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K. Again we have opted to show only one direction (the same direction as in Figure 6(a)) for simplicity. All compounds near saturation by 7 T, though Er2-xTi6+xBi9 has not saturated, owing to an unusual inflection in the data near 4 T. Regardless, we can extract the approximate saturation magnetization for all compounds, which has been summarized in Table 1 (column M). Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2-xTi6+xBi9. This will underestimate the true saturation magnetization, as we will be normalizing to 2 rare-earth atoms per formula unit instead of (2-x). However, we can compare these results with the expected saturation magnetization from XRD and EDS results (M and M). Comparing directly, we find that the experimental results actually agree quite well with the SCXRD results. Some other parameters of interest (e.g. and ) have also been added to Table 1. Due to the disorder and variable number of Ln sites, we consider these values as first approximations.
Despite the high levels of disorder, some samples of Ln2-xTi6+xBi9 (e.g. Tb) appear to show magnetization consistent with long-range order. This is highly unexpected, and considering the lack of ZFC/FC splitting and no clear frequency dependence, more detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed.
IV Conclusions
We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, a kagome metal belonging to the AM3X4 family of materials. We described the rational for the formation of TbTi3Bi4, including the increased competition with another class of materials, the Ln2-xTi6+xBi9 shurikagome metals for small rare-earth species (Ln: Tb–Lu). Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases surrounding multiple metamagnetic transitions. Transport measurements highlight the impact the rare-earth magnetism has on the kagome-dominated transport. Upon entering the magnetization plateau, the system sharply switches between positive (30-40%) and negative (-55%) magnetoresistance in a system that is already quite metallic (1 Ohm-cm). We have also briefly examined the properties of the Ln2-xTi6+xBi9 shurikagome metals. Despite considerable disorder on the rare-earth sublattices, some systems exhibit magnetization responses more indicative of long-range order. Together, these results not only provide insight into a new kagome lattice and an unusual family of disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
V Acknowledgements
The work of B.R.O (synthesis, data collection, writing), H.Z. (analysis), K.G. (He3 data collection), A.F.M. (He3 data collection), and M.A.M. (editing), was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division. We thank the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source for use of their Photonics Laue Camera System.

### v2

@@
-Complex magnetic landscape in kagome metal TbTi3Bi4 and competition with family of Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals
+Intricate magnetic landscape in antiferromagnetic kagome metal TbTi3Bi4 and interplay with Ln2-xTi6+xBi9 (Ln: TbLu) shurikagome metals
 Abstract
-Here we present the discovery and characterization of TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 family.
-We previously reported on the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) kagome metals.
-This structure type represents a magnetically diverse and exfoliable family of kagome metals.
-Detailed characterization of the magnetic and electronic transport properties of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state.
-The system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (-50%) near a metamagnetic transition.
-Unlike the Lu-Gd analogs, TbTi3Bi4 cannot be synthesized by the methodology presented previously due to phase competition with another, unreported class of materials, the Ln2-xTi6+xBi9 (x1.7–1.2) (Ln: Tb–Lu) shurikagome metals.
+Here we present the discovery and characterization of the kagome metal TbTi3Bi4 in tandem with a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals.
+We previously reported on the growth of the LnTi3Bi4 (Ln: La–Gd3+, Eu2+, Yb2+) family, a chemically diverse and exfoliable series of kagome metals with complex and highly anisotropic magnetism.
+However, unlike the La–Gd analogs, TbTi3Bi4 cannot be synthesized by our previous methodology due to phase competition with Ln2-xTi6+xBi9 (x1.7–1.2).
+Here we discuss the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families, helping to frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and providing a strategy to circumvent formation of Ln2-xTi6+xBi9.
+Detailed characterization of the magnetic and electronic transport properties on single crystals of TbTi3Bi4 reveals a highly complex landscape of magnetic phases arising from an antiferromagnetic ground state.
+A series of metamagnetic transitions creates at least 5 unique magnetic phase pockets, including a 1/3 and 2/3 magnetization plateau.
+Further, the system exhibits an intimate connection between the magnetism and magnetotransport, exhibiting sharp switching from positive (+40%) to negative magnetoresistance (–50%).
 Like the LnTi3Bi4 kagome metals, the Ln2-xTi6+xBi9 family exhibits quasi-2D networks of titanium and chains of rare-earth.
-We present the structures and some select basic properties of the Ln2-xTi6+xBi9 family alongside the newly discovered TbTi3Bi4.
-A discussion regarding the phase competition between the LnTi3Bi4 and Ln2-xTi6+xBi9 families helps frame the difficulty in synthesizing LnTi3Bi4 compounds with small Ln species and provides a strategy to circumvent formation of Ln2-xTi6+xBi9.
+We present the structures and some basic magnetic properties of the Ln2-xTi6+xBi9 family alongside our characterization of the newly discovered TbTi3Bi4.
 111Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE).
@@
 Spurred on by continued discoveries in the AM3X5, AM6X6, and CoSn-type compounds, kagome metals have proven to be a rich space for exploring the intersection of solid-state chemistry and fundamental condensed matter physics.
-The impetus for much of the inspiration comes from the basic electronic structure afforded by the kagome motif – one which can contain Dirac points, flat bands, and Van Hove singularities.
+The impetus for studying kagome metals is derived from the basic electronic structure of a kagome toy model – which creates the potential for Dirac points, flat bands, and Van Hove singularities.
 [1, 2, 3, 4].
-Derived from the triangular lattice, the kagome motif have shown remarkable structural and chemical diversity, allowing for a rich interaction between structural chemistry, electronic properties, and magnetism.
-As a prime example, the AM3X5 compounds have been heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, superconductivity).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity.[14] In fact, embedding magnetism into the kagome motif underpins much of the research into the AM6X6 and CoSn-type compounds, which continues to provide the motivation for the development of new kagome materials.[15, 16, 17, 18, 16, 19, 20, 21, 22, 23, 24, 25, 26]
-One fundamental way to develop new materials is to perturb simpler, “fundamental” prototype structures.
-Take, for example, the CoSn family of compounds.
-By adding a filler atom in the natural voids formed by the structure, the CoSn-type stacking gives way to another of the most researched kagome families, the AM6X6.
-Further perturbations on the AM6X6 prototype yield a grandchild of the CoSn prototype – the AM3X4 kagome metals.
-We reported previously on the discovery of both the antimonide AV3Sb4 (A: Eu2+, Yb2+) [27] and bismuthide ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[28] which make up a larger family of AM3X4 compounds.[27, 28, 29, 30, 31, 32, 33] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals.
+Indeed, the latent potential within the kagome motif is often unlocked by designing materials where the Fermi level (E) is aligned with these electronic features, as exemplified by the AV3Sb5 (A: K, Rb, Cs) kagome materials.
+In AV3Sb5 compounds, the Fermi level is intrinsically tuned to a Van Hove singularity, causing these nonmagnetic materials to be heavily investigated for complex interactions between electronic and structural instabilities (charge density wave, potential unconventional superconductivity, non-trivial topology).[5, 6, 7, 8, 9, 10, 11, 12, 13] The recent discovery of CsCr3Sb5 has also included the possibility of exploring the interaction between a possible spin density wave and superconductivity in proximity to a flat band near E.[14]
+Indeed, CsCr3Sb5 exemplifies another key direction for the exploration of new kagome materials – the interaction between magnetic sublattices and the underlying kagome band structure.
+The development of new magnetic kagome systems (either with magnetic kagome sublattices, or magnetic supporting sublattices) has been remarkable.
+Consider, for example, the chemical diversity of the CoSn-family and it’s grandchildren, the AM6X6 compounds.[15, 16, 17, 18, 19, 17, 20, 21, 22, 23, 24, 25, 26, 27] The structural relationship between CoSn and the AM6X6 is an instructive one, as the two compounds are fundamentally linked.
+The CoSn prototype features large interstitial voids between the hexagonal plaquettes of adjacent kagome networks, which offers a unique opportunity for generating new kagome materials.
+By adding an electropositive filler atom in these natural voids, we can produce ordered superlattices – giving birth to the AM6X6 (HfFe6Sn6, P6/mmm) prototype.[28] The more complex ordering of the AM6X6 creates an explosion of structural diversity, as the ordered superlattice exhibits multiple stacking variations and intergrowths like ErFe6Sn6 (Amam), DyFe6Sn6 (Cmmm), and YFe6Sn6 (Immm).[29] The titular series – the AM3X4 kagome metals – also contain remnants of the AM6X6 stacking, producing a prototypical grandchild of the CoSn family.
+We previously reported the discovery of both the antimonide LnV3Sb4 (Ln: Eu2+, Yb2+) [30] and bismuthide LnTi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals,[31] which make up a larger family of AM3X4 compounds.[32, 31, 33, 34, 35, 36, 37] This family boasts both magnetic and non-magnetic compounds through the diverse choice of A cations, which are typically electropositive rare-earth and alkaline-earth metals.
 The A-site elements comprise zig-zag chains which run parallel to the M-site kagome layers.
-Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds maintain a much more quasi-2D crystal structure.
-Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure.
-Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission (ARPES).[32, 34, 35, 36, 37, 38]
+Remarkably, unlike the CoSn and AM6X6 parent structures, the AM3X4 compounds exhibit a much more quasi-2D, layered, exfoliable crystal structure.
+Developing quasi-2D kagome platforms is critical, as the dimensionality of the kagome network and the supporting sublattices have a profound effect on the realization of key features in the electronic structure.[38] Crystals are easily exfoliated (similar to the AM3X5 family), which makes them an attractive option for exploring both magnetism and the kagome band structure.
+Accordingly, the AM3X4 compounds have been the focus of a growing body of work, particularly studies involving angle-resolved photoemission spectroscopy (ARPES).[36, 39, 40, 41, 42, 43]
 Under the conditions presented in our previous work, LnTi3Bi4 compounds containing rare-earth atoms smaller than Gd3+ did not appear.
 In this work, we present a modified growth procedure which allows stabilization of TbTi3Bi4, the smallest rare-earth cation reported in the LnTi3Bi4 structure thus far.
-Our investigation into the smaller rare-earth species (Tb–Lu) also unveiled an intimate competition between the LnTi3Bi4 kagome metals and a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals.
-By understanding the phase diagram surrounding the LnTi3Bi4 compound, we are able to subvert the formation of Tb2-xTi6+xBi9 and stabilize single crystals of TbTi3Bi4.
+Our work suggests that TbTi3Bi4 is a relatively low temperature phase (500°C) and is suppressed through competition with a new series of compounds, the Ln2-xTi6+xBi9 (Ln: Tb–Lu) shurikagome metals.
+By quenching fluxes from 1100°C to 600°C, we suppress the growth of Tb2-xTi6+xBi9 and encourage the resorption of any shurikagome nuclei.
+A subsequent cool from 600°C to 400°C encourages the growth of single crystals of TbTi3Bi4.
 This work details our rationale for the stabilization of TbTi3Bi4 alongside a discussion of its magnetic and electronic transport properties.
 Remarkably, TbTi3Bi4 presents some of the most rich and complex magnetism in the known AM3X4 compounds, reinforcing the need to understand and control the formation of new phases.
-Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals, which have strong chemical disorder and unusual magnetic behavior.
+Our work also briefly investigates the structural and magnetic properties of the new Ln2-xTi6+xBi9 (Ln: Tb–Lu) metals.
 II Methods
@@
 TbTi3Bi4 single crystals are grown through a bismuth self-flux.
-Tb metal (Ames), Ti (Alfa 99.9% powder), and Bi (Alfa 99.999% low-oxide shot) were measured in a 2:3:20 ratio into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit.[39] The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules.
-The flux was heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h.
-The sample was then quenched by removing it from the furnace and immediately placing in a preheated 600°C furnace.
-The growth proceeded with a cooling rate of 2°C/hr to 400°C, where excess bismuth was removed through centrifugation.
-Ln2-xTi6+xBi9 single crystals are grown in a similar manner, except the samples are cooled linearly from 1100°C to 600°C at a rate of 2°C/hr before centrifugation at 600°C.
+Tb metal (Ames), Ti powder (Alfa 99.9%), and Bi shot (Alfa 99.999% low-oxide) were placed into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit[44] at a ratio of 2:3:20 Tb:Ti:Bi.
+The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules.
+Samples were subsequently heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h.
+The samples were then quenched by removing them from the furnace and immediately placing them into a different furnace preheated to 600°C.
+The growth proceeded with a cooling rate of 1°C/hr to 400°C.
+Excess bismuth was removed through centrifugation at 400°C.
+Ln2-xTi6+xBi9 single crystals are grown through a bismuth self-flux as well.
+Ln metal (Ames), Ti powder (Alfa 99.9%), and Bi shot (Alfa 99.999% low-oxide) were placed into a 5 mL Canfield crucibles fitted with a catch crucible and a porous frit[44] at a ratio of 2:3:20 Ln:Ti:Bi.
+The crucibles were sealed under approximately 0.7 atm of argon gas in fused silica ampoules.
+Ln2-xTi6+xBi9 single crystals are stable over a wide range of fluxes (e.g., Ln:Ti:Bi of 2:3:12, 2:3:20, 1:1:3).
+Crystals featured here were grown at a ratio of 2:3:12, though the composition of the crystals did not change appreciably under the various flux compositions investigated here.
+In particular, we are currently unable to grow the idealized, fully saturated Ln2Ti6Bi9.
+Fluxes were heated to 1100°C at a rate of 200°C/hr and thermalized at 1100°C for 12 h.
+The samples were then cooled linearly from 1100°C to 600°C at a rate of 2°C/hr.
+Excess bismuth was removed through centrifugation at 600°C.
 II.2 Bulk Characterization
-Single crystals were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD).
+Single crystals of both TbTi3Bi4 and Ln2-xTi6+xBi9 were mounted on kapton loops with Paratone oil for single crystal x-ray diffraction (SCXRD).
 Diffraction data were collected at 100 K on a Bruker D8 Advance Quest diffractometer with a graphite monochromator using Mo K radiation ( = 0.71073 Å).
@@
 CIF files for all structures are included in the ESI.
-Laue diffraction utilizing a Photonic Science vertical Laue system was used to orient crystals of TbTi3Bi4 for physical property measurements.
-Magnetization measurements of LnTi3Bi4 single crystals were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode.
-Additional measurements below 1.8 K on the MPMS3 utilized an additional He3 insert.
-Angle-resolved magnetization measurements were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
-Heat capacity measurements were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS) equipped a dilution refrigerator insert.
+A TF20 X-Ray System (LAUE-V-674-AC-GigE from Photonic Science) was used to characterize and orient single crystals for magnetotransport and magnetization measurements.
+Elemental analysis was carried out on as-grown crystals using a Hitachi-TM3000 microscope equipped with a Bruker Quantax 70 EDS system.
+Magnetization measurements (300 K – 1.8 K) on crystals of TbTi3Bi4 and Ln2-xTi6+xBi9 were performed in a 7 T Quantum Design Magnetic Property Measurement System (MPMS3) SQUID magnetometer in vibrating-sample magnetometry (VSM) mode.
+Additional measurements on TbTi3Bi4 (1.8 K – 0.3 K) utilized the Quantum Design iHe-3 He3 insert for the MPMS3.
+AC susceptibility measurements (3.5 K – 0.1 K) on Tm2-xTi6+xBi9 utilized a Quantum Design Dilution Refrigerator insert in a Quantum Design 9 T Dynacool PPMS.
+The lowest frequency data (100 Hz) was normalized to the MPMS3 data from 4 K to 1.8 K to create a continuous data set from 300 K to 0.1 K.
+Angle-resolved magnetization measurements on TbTi3Bi4 were performed on a 7 T Quantum Design Magnetic Property Measurement System (MPMSXL) equipped with a rotator stage.
+A note regarding hysteresis in TbTi3Bi4; crystals show substantial magnetic hysteresis, and considerable care must be taken to account for magnetic and thermal history.
+While samples show negligible changes in low-field (500 Oe) measurements under field-cooled (FC) or zero-field-cooled (ZFC) conditions, field loops up to 3 T must be performed with consideration for the ZFC (virgin magnetization) curve.
+Additionally, subsequent measurements under varying temperatures (e.g.
+M vs.
+H curves at different temperatures) need to heat up above the transition temperature to “erase” the magnetic history.
+Specific details for each magnetic and transport measurement are included in the text alongside the respective figures.
+Electronic transport measurements on TbTi3Bi4 were performed in a Quantum Design 9 T Dynacool PPMS.
+Single crystals were mounted to a sheet of sapphire, which was subsequently adhered to the sample puck stage using GE varish.
+The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane.
+Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%) in a six-wire geometry for simultaneous measurements of the Hall effect and magnetoresistance.
+As alluded to in the magnetization methods, care must be taken to account for magnetic and thermal history.
+We are primarily focused on the magnetotransport results that are repeatable under continued field cycling, and will typically omit virgin magnetoresistance curves.
+In this case, we typically begin each measurement in a fully-polarized (e.g.
+9 T) state.
+All transport measurements utilized a probe current of 8 mA.
+Heat capacity measurements (300–1.8 K) on TbTi3Bi4, Tb2-xTi6+xBi9, and Er2-xTi6+xBi9 were performed in a Quantum Design 9 T Dynacool Physical Property Measurement System (PPMS).
+Additional heat capacity measurements (3.5–0.1 K) on TbTi3Bi4 utilized the Quantum Design Dilution Refrigerator insert for the Quantum Design 9 T Dynacool PPMS.
 Thermal contact was achieved using Apezion N-grease.
-Electronic resistivity measurements were performed in a Quantum Design 9 T Dynacool PPMS.
-Single crystals were mounted to a sheet of sapphire and then the sapphire plate was then adhered to the sample puck stage using GE varish.
-The sapphire could be rotated between different orientations to achieve measurements with the magnetic field both within and out-of-plane.
-Sample surfaces were cleaned by removing several layers using adhesive tape, and then contacts were made using silver paint (DuPont cp4929N-100) and platinum wire (Alfa, 0.05 mm Premion 99.995%).
+II.3 Electronic Structure Calculations
+In an effort to describe the underlying electronic structure of the Ln2Ti6Bi9 family, we have performed first-principles all-electron density functional theory calculations within the generalized gradient approximation, as in several of our recent works.[31, 45, 46, 47, 48] To avoid 4f-electron-related complications we have chosen to study the Lu-based end-member Lu2Ti6Bi9 in a non-spin-polarized configuration, with Lu having a full 4f shell and thereby no magnetic character.
+We have chosen to neglect potential structural disorder by studying the fully-ordered, stoichiometric structure of Lu2Ti6Bi9.
+The results should be generally applicable to other lanthanides in this family, presuming that relevant measurements (such as ARPES) are performed at temperatures well above the onset of any magnetic order.[49] This is as in our previous work.[31] The linearized augmented plane-wave (LAPW) basis was employed, with an RKmax (the product of the smallest muffin-tin sphere radius and largest planewave expansion wavevector) of 8.0 and sphere radii of 2.50 Bohr for Bi and Lu were chosen, and 2.41 Bohr for Ti.
+Neither spin-orbit coupling nor the GGA+U approach (i.e., a ‘U’) was included.
 III Results and Discussion
 III.1 Crystal Structure and Synthesis
-Our prior work identified that the ATi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[28] However, growth of ATi3Bi4 crystals containing smaller rare-earth metals (Tb–Lu) was stymied due to preferential formation of Ln2-xTi6+xBi9.
+Our prior work identified that the LnTi3Bi4 (Ln: La…Gd3+, Eu2+, Yb2+) kagome metals are readily grown from a bismuth self-flux.[31] However, growth of LnTi3Bi4 crystals containing smaller rare-earth metals (Ln: Tb–Lu) was stymied due to extensive crystallization of the Ln2-xTi6+xBi9 compounds.
 Figure 1(a) provides a simple schematic showing the lanthanide row sorted in order of the 9-coordinate Shannon ionic radius.
-Within the ATi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state.
+Within the LnTi3Bi4 compounds, both Eu and Yb appear to preferentially adopt the 2+ oxidation state.
 As such, the radii of Eu2+ and Yb2+ are analogous to the larger La3+ and Pr3+, respectively.
-Large cations preferentially stabilize the ATi3Bi4 phase (denoted by blue), and small cations preferentially stabilize Ln2-xTi6+xBi9 (green).
+Large cations preferentially stabilize the LnTi3Bi4 phase, and small cations preferentially stabilize Ln2-xTi6+xBi9.
 Within this work, we demonstrate that Tb represents a unique cross-over between the two phases, and both TbTi3Bi4 and Tb2-xTi6+xBi9 can be obtained by tuning the growth profiles.
-Figure 1(b,c) details the crystal structure of both the ATi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds.
-As we have discussed ATi3Bi4 extensively before, for simplicity we opt to show only half the unit cell here (the omitted half is generated through an inversion operation, space group Fmmm.
-ATi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln-Ln zig-zag chains.
-The local coordination of the Ln species is approximately 9-fold.
-Crystals of ATi3Bi4 are quasi-2D and can be easily exfoliated using adhesive tape.
-We have observed that both M-Sb and Ln-Sb terminations can be achieved.[27] The kagome network in ATi3Bi4 compounds is only slightly distorted, though the quasi-1D zig-zag chain limits the symmetry of the system to orthorhombic.
-The Ln2-xTi6+xBi9 compounds can be understood as a heavily alloyed form of Ti8Bi9[40], where Ln atoms preferentially occupy two of the Ti sites.
+Figure 1(b,c) details the crystal structure of both the LnTi3Bi4 (b) and Ln2-xTi6+xBi9 (c) compounds.
+Note that Figure 1(b) presents only half of the LnTi3Bi4 unit cell, with the omitted half generated by an inversion operation (space group Fmmm).
+LnTi3Bi4 contains two key structural motifs of interest; 1) a Ti–Ti kagome network, and 2) Ln–Ln zig-zag chains.
+The reduced symmetry of the unit cell is driven by the quasi-1D zig-zag chains, which necessitate a maximum of 2-fold rotational symmetry.
+Despite the orthorhombic structure, the kagome network is negligibly distorted (0.1Å out-of-plane buckling).
+Accordingly, the orthorhombic b/a lattice parameter ratio () is very nearly the nominal applicable to the hexagonal symmetry.
+The zig-zag chains dominate the magnetic properties of these materials, with the nearest-neighbor intra-chain Ln–Ln distance (4 Å) substantially closer than the inter-chain distances (6 Å).
+We typically model the rare-earth local coordination as approximately 9-fold, though the coordination is not uniform and requires a small range of bond lengths (3.3 Å– 3.6 Å) to recover the polyhedral environment shown in Figure 1(b).
+The Ln2-xTi6+xBi9 (P4/nmm) compounds can be understood as a heavily alloyed form of Ti8Bi9[50], where Ln atoms preferentially occupy two of the Ti sites (there are 4 unique Ti sites in Ti8Bi9).
 If the substitution was perfect, (e.g.
 Ln2Ti6Bi9) the Ln atoms would create a nearly ideal 1D chain that cuts perpendicular to the Ti–Ti network.
-The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to ATi3Bi4 (4Å)The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles.
-Paying homage to both the linguistic origin and coordination of the “kagome lattice,” the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
-The two sites that are preferentially occupied by Ln atoms possess considerably different coordination than the remaining Ti sites in the parent Ti8Bi9 structure.
+The Ln–Ln distance is closer in Ln2Ti6Bi9 (3.8Å) when compared to LnTi3Bi4 (4Å), owing to the contraction from partial occupancy of Ti.
+The remaining Ti atoms form a unique network made of up Ti4 tetrahedra interconnected with corner-sharing triangles.
+Paying homage to both the linguistic origin and the corner-sharing coordination of the kagome lattice, the tessellated unit resembles a four-pointed hira-shuriken and we’ve taken to nicknaming these compounds “shurikagome” metals.
+Within Ti8Bi9, the two sites that are preferentially occupied by Ln in Ln2Ti6Bi9 (Ln4 and Ln5 in Figure 1(c)) possess considerably different coordination than the Ti that comprise the shurikagome network.
 Considering bonds 3.2 Å, Ti atoms within the shurikagome network have an approximate coordination of 10 (4 Ti–Ti bonds, 6 Ti–Bi bonds) and a polyhedral volume of approximately 50.7 Å3.
-Sites preferentially occupied by Ln possess a coordination of 8, solely from Ln–Bi bonds, and form slightly distorted square anti-prisms.
-The two mixed Ln/Ti sites are also distinct from one another, and one has a substantially larger polyhedral volume (56.9Å3) than the other (49.5Å3).
-As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and further preferentially occupy the largest of the mixed Ln/Ti sites.
-Generally speaking, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site.
+However, the Ln4 and Ln5 sites possess a coordination of 8, entirely from Ln–Bi bonds, and form slightly distorted square anti-prisms.
+Ln4 and Ln5 are also distinct from each other.
+Ln4 sits within the plane of the shurikagome network, and has substantially larger polyhedral volume (56.9 Å3) owing to the support of the Ti–Ti scaffolding.
+The Ln5 site is slightly collapsed along c, leading to a correspondingly lower volume (49.5 Å3) and a marginally more distorted coordination.
+As a conceptually pleasing observation, the rare-earth atoms appear to exclusively occupy the 8-coordinate sites, and preferentially occupy the largest of the mixed Ln/Ti sites.
+From single crystal diffraction results, the larger site is approximately 75–90% occupied by rare-earth, compared to 50–75% for the smaller site.
 As such, the total stoichiometry can be more precisely written as Ln2-xTi6+xBi9 (x1.7–1.2).
-CIF files for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI.
+CIF files and associated crystallograpic tables for both TbTi3Bi4 and the Ln2-xTi6+xBi9 compounds are included in the ESI.
 We will discuss trends in the Ln substitution alongside the magnetic properties in a later section.
-As alluded to before, TbTi3Bi4 is not stable under conditions presented in our first manuscript.[28] Samples of large rare-earth ATi3Bi4 compounds are readily grown through a simple flux (cooled from 1100°C to 600°C at a rate of 2°C/hr).
-However, samples containing Tb will exclusively form Tb2-xTi6+xBi9.
+As alluded to before, TbTi3Bi4 does not appear to form under conditions presented in our first manuscript.[31] Samples of large rare-earth LnTi3Bi4 compounds are readily grown through a bismuth self-flux cooled from 1100°C to 600°C.
+However, reactions utilizing Tb will exclusively form Tb2-xTi6+xBi9.
+To help understand the shift towards crystallization of the Tb2-xTi6+xBi9 phase, we first turned to examine the known phases in the vicinity of TbTi3Bi4.
 Figure 2(a) shows a small section of the Tb–Ti–Bi phase diagram near several of the phases of interest.
-Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[41, 42] Note the proximity of TbTi3Bi4 and Tb2-xTi6+xBi9 (particularly the proximity of the the hypothetically ideal Tb2Ti6Bi9).
-The large single-phase region of Ti8Bi9 (and the associated alloys with Tb2-xTi6+xBi9) is problematic, as it isolates a large swath of the phase diagram from Bi-rich fluxes.
-The single phase region even persists beyond where a hypothetical Bi2Ti6Bi9 exists (complete substitution of Bi on the mixed sites).
-There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[43] However, if Tb2Ti6Bi9 is not stable, then TbTi3Sb4 should still be accessible from a Bi self-flux.
-We designed a series of experiments to test both the thermal stability of Tb2-xTi6+xBi9 and compositional dependence.
-We first observed that samples allowed to completely freeze (T=270°C) often exhibited Tb2-xTi6+xBi9 crystals that were rounded, as though they started to decompose.
-Modifying the flux towards Tb-rich conditions instead resulted in rounded crystals of TbBi.
-In both cases, the corners of the decomposing crystals were encased in agglomerates of TbTi3Bi4.
-This highlighted that TbTi3Bi4 is a low-temperature phase.
+Data for the existing binary and pseudobinary lines, including the single phase region surrounding Ti8Bi9 were adapted from published Ti–Bi, Tb–Bi binary diagrams.[51, 52] The large Ti–Bi offstoichiometry in samples of Ti8Bi9 is remarkable, and suggests that solid solubility exits beyond where the hypothetical compound Bi2Ti6Bi9 exists (written in this form to highlight substitution of Bi on the mixed sites).
+There have also been reports that oxygen atoms can substitute in the center of the Ti4 tetrahedra, creating compositions like Ti8Bi9O0.25.[53] Such observations are consistent with the structural flexibility needed to host the Ln disorder proposed here.
+It appears that the fully-occupied Tb2Ti6Bi9 is not stable under conditions explored here, as flux growths containing different concentrations of Ln:Ti:Bi consistently produce samples of Tb2-xTi6+xBi9 which are sub-occupied.
+This is fortunate from the perspective of TbTi3Bi4, as the presence of Tb2Ti6Bi9 would hinder our ability to form TbTi3Bi4 from a bismuth-rich (low-temperature melting point) mixture.
+Given the absence of Tb2Ti6Bi9, we hypothesized that TbTi3Bi4 could potentially be accessible from a Bi self-flux.
+However, if the phase exists, it must be a low-temperature phase, which is out-competed by the thermal stability of Tb2-xTi6+xBi9.
+One preliminary observation that supported our hypothesis was the rounding (resorption) of the corners and edges of Tb2-xTi6+xBi9 crystals at temperatures near 600°C (flux composition 2:3:20).
+Samples quenched at higher temperatures (e.g.
+800°C) showed nicely faceted Tb2-xTi6+xBi9 samples with no resorption, suggesting that the shurikagome is either: 1) unstable at low temperatures, or 2) is consumed by crystallization of a low-temperature phase.
+A similar phenomenon was noted with more Tb-rich compositions (flux composition 5:5:20) and the high melting point TbBi binary.
+Given sufficient time at low-temperatures, the decomposed corners of Tb2-xTi6+xBi9 became encased in recrystallized agglomerates of TbTi3Bi4.
+This highlighted that TbTi3Bi4 is a low-temperature phase and evolves from the dissolution of the high-temperature compounds.
 Subsequent quenching experiments further suggested that TbTi3Bi4 is likely only stable below 600°C.
-Samples grown above 600°C reveal two alkemade triangles; 1) three phase TbBi–Tb2-xTi6+xBi9–Bi and 2) two phase Tb2-xTi6+xBi9–Bi.
-Samples grown below 600°C reveal instead; 1) three phase TbBi–TbTi3Bi4–Bi and 2) three phase TbTi3Bi4–Tb2-xTi6+xBi9–Bi.
-Figure 2(b) presents a simple projection of Figure 2(a) along the TiBi–TbBi pseudobinary line, summarizing the relative proximity and temperature stability of these results.
-Crystals of TbTi3Sb4 are 1-5 mm pseudo-hexagonal crystals with a brilliant metallic luster.
-They are readily exfoliated and are lightly air-sensitive.
-Samples can be worked with in air, though they will swell and spall under extended (1 day) exposure.
-Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster.
-Samples are readily cleaved with a blade parallel to the ab-plane.
+Thus, we propose that TbTi3Bi4 is a low-temperature phase with a maximum stability of approximately 600°C in a bismuth self-flux.
+Upon cooling, it is preceded by either TbBi or Tb2-xTi6+xBi9, depending on the flux composition.
+To circumvent formation of the competing phases, we quench samples from 1100° C to 600°C.
+This step prevents extended nucleation and growth of Tb2-xTi6+xBi9.
+The samples are then slow-cooled from 600°C to 400°C at a rate of 1°C/hr, ensuring resorption of any shurikagome nuclei and encouraging growth of TbTi3Bi4.
+Batches grown in this manner are primarily composed of nicely-faceted TbTi3Bi4.
+We are pleased to see that these observations agree nicely with recent results from other groups.[54, 55]
+Crystals of TbTi3Bi4 are 1–5 mm pseudo-hexagonal crystals with a brilliant metallic luster.
+They are extremely soft, readily exfoliated parallel to the ab plane, and are lightly air-sensitive.
+In the antimonide analog, we have observed that both M-Sb and Ln-Sb terminations can be achieved through exfoliation, and we expect similar results for the bismuthide.[32] Crystals of Ln2-xTi6+xBi9 are typically truncated square plates (nearly octagonal) with a dark gray luster.
+Crystals are substantially more brittle than the LnTi3Bi4 compounds but can be readily cleaved parallel to the ab-plane (perpendicular to the Ln chains) with a sharp blade.
 Ln2-xTi6+xBi9 crystals are substantially more sensitive to air, and crystals will completely disintegrate within a few hours.
-In both cases, the samples expand and self-cleave along the ab-plane upon extended exposure to air.
+Curiously, both samples decompose in the same manner.
+Upon extended exposure, the samples swell substantially and spall layer-by-layer.
 III.2 Properties of TbTi3Bi4
-Due to the reduced symmetry (Fmmm), the ATi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains.
+Due to the reduced symmetry (Fmmm), the LnTi3Bi4 materials tend to be highly anisotropic, particularly with regards to the quasi-1D zig-zag chains.
 TbTi3Bi4 is no exception.
 To diagnose the potential anisotropies in the material, we first begin with a study of the magnetization along different principle crystallograpic axes.
-Figure 3(a) presents the magnetization of a single crystal of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions.
-TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two primary transitions at 24 K and 3 K.
+Figure 3(a) presents the field-cooled (FC) magnetization curve of a single crystal of TbTi3Bi4 with the magnetic field (100 Oe) parallel to the [100], [010], and [001] directions.
+At these low fields, there is no difference between the field-cooled (FC) and zero-field-cooled (ZFC) measurements.
+TbTi3Bi4 appears to adopt an antiferromagnetic ground state, with two transitions at 24 K and 3 K.
 The [100] direction corresponds to the “easy axis,” and orienting the magnetic field parallel to the [010] or [001] induces a substantially weaker effect.
-The behavior is reminiscent of GdTi3Bi4[28], though crystals of TbTi3Bi4 respond negligibly to fields oriented along [001].
-Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions.
-Upon cursory inspection, the data again resembled that of GdTi3Bi4 (two metamagnetic transitions when ).
-However, the metamagnetic transitions in TbTi3Bi4 exhibit substantial field hysteresis.
-The data saturates at approximately The first transition plateaus at approximately 9.2 , in good agreement with the 9 expected for Tb3+.
-The first metamagnetic transition first plateaus at a moment of 2.9-3.2 , which is approximately 1/3 of the expected net moment.
-Furthermore, closer inspection at the second metamagnetic transition reveals an unusual inflection midway through the rising field direction (see Figure 3(b,inset)).
-This inflection occurs at approximately 6.2 slightly above 2/3 the expected moment.
-This inflection and other properties of the isothermal magnetization curves will be studied in detail below.
-Similar to the temperature-dependent magnetization, there are only weak, linear responses when or .
+Figure 3(b) demonstrates the isothermal magnetization of TbTi3Bi4 with the magnetic field parallel to the [100], [010], and [001] directions at 1.8 K.
+Due to the hysteretic nature of TbTi3Bi4, care must be taken to control the magnetic and thermal history of the samples.
+For example, Figure 3(b) () shows three different field paths: 1) ZFC virgin magnetization data (gray), 2) decreasing field (purple, down arrow), and 3) increasing field (purple, up arrow).
+The virgin curve is only obtained upon the first field sweep after cooling below 24 K with no applied field (ZFC).
+All subsequent field sweeps will fall onto the purple data, tracing either the field increasing (up arrow) or field decreasing (down arrow) path.
+Continuing the data into the negative field direction is not necessary as the system exhibits no coercivity at zero-field, and loops collected from 0 T to 7 T are functionally identical to those from 0 T to 7 T.
+For the vast majority of our analysis we will omit the virgin magnetization curves.
+The differences between the ZFC virgin curve and the next positive field sweep are likely due to the domain structure that forms on cooling in zero field, and are not the focus of this work.
+The isothermal magnetization data saturates at approximately 9.2 by 3 T, in good agreement with the M expected for Tb3+.
+The first metamagnetic transition plateaus at a moment of 2.9 –3.2 , which is approximately 1/3 M.
+Closer inspection midway through the second metamagnetic transition reveals an unusual inflection (see Figure 3(b,inset)).
+This inflection occurs at approximately 6.2 , slightly above 2/3 M.
+This inflection and other properties of the isothermal magnetization curves will be studied in the following figure.
+As implied by the temperature-dependent magnetization shown in Figure 3(a), the magnetic response is suppressed quickly as we rotate the field away from the [100] direction.
 Figure 3(c) provides a series of magnetization curves collected while rotating between , , and in 15° increments.
-The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further easy-[100]) anisotropy.
-Figure 3(c) presents the inverse susceptibility, collected under a 100 Oe field where .
+Only the increasing and decreasing field directions are shown (no ZFC virgin magnetization curve).
+The angular dependence is relatively simple, with the transitions being stretched towards higher fields with increasing deviations from , again suggesting strong easy-plane (and further, easy-[100]) anisotropy.
+Figure 3(d) presents the inverse susceptibility, collected under a 100 Oe (FC) field where .
 The high temperature data exhibits a slight upwards curvature, potentially due to a temperature-independent contribution from core diamagnetism (bismuth) or the quartz sample holder.
-We show Curie-Weiss fits to the data with (gray) and without (black) a temperature-independent term.
-While the gray curve fits over a longer range of data, it requires a large emu mol-1 Oe-1 and yields an unphysical effective moment of 11 .
-We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit, which yields a =53 K and a = 9.9.
-The effective moment agrees well with that expected of Tb3+ (9.72), though the Weiss temperature is unphysically large (53 K), likely affected by the temperature independent contribution, limited temperature range, and crystal-field effects.
-As a consistency check with the magnetization data, we performed zero-field heat capacity measurements down to 100 mK.
-The sample was initially oriented with such that measurements could also be made for the intermediate plateaus seen in Figure 2(b).
-The first order hysteretic nature of the transitions seen in the isothermal magnetization made the quantative analysis of field-dependent heat capacity challenging, and we have opted to reserve this data for a later study.
-Instead, we focus on the analysis of the zero-field heat capacity utilizing samples of LaTi3Bi4 as a nonmagnetic standard.
-Figure 3(d) demonstrates the experimental heat capacity data (green) as a function of temperature, plotted in .
+To account for the curvature, we attempted fits utilizing a temperature-independent term, though the result was an nonphysical effective moment of 11 .
+We opt instead to utilize a limited high-temperature regime (red) for a simpler Curie-Weiss fit (black), which yields a =53 K and a = 9.9.
+The effective moment agrees well with that expected of Tb3+ (9.72).
+However, the Weiss temperature is unusually large (53 K), likely affected by the temperature independent contribution, limited temperature range of the fit, and crystal-field effects.
+Figure 3(e) presents the zero-field heat capacity (plotted in ) data collected on a single crystal of TbTi3Bi4 down to 100 mK with measurements of LaTi3Bi4 used as a nonmagnetic standard.
+Two data sets were combined to construct this data set, a conventional helium measurement (200 K – 1.8 K) and a measurement performed with a dilution refrigerator insert (4 K – 100 mK).
+The data sets did not require substantial adjustment or scaling to overlap in the region from 4 K – 1.8 K after normalization by the sample mass.
 The two magnetic transitions are evident around 20 K and 2 K.
@@
 The peak is still increasing at 0.2 K, which agrees relatively well with a Tb-based Schottky peak ( K).[56] Utilizing LaTi3Bi4 as the nonmagnetic standard (black) and invoking a simple nuclear Schottky model , we can extract the magnetic entropy for TbTi3Bi4.
-Figure 3(e) shows the integrated magnetic entropy (green) as a function of temperature.
+Figure 3(f) shows the integrated magnetic entropy (green) as a function of temperature.
 The extracted magnetic contribution to the heat capacity () is also shown (gray) as a schematic reference for the magnetic transition temperatures (not to scale).
 Interestingly, integrating up to the primary magnetic transition appears to correspond with an entropy release of approximately .
-Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other ATi3Bi4 materials.[28] By 100 K we recover approximately 90% of the total entropy expected for Tb3+.
-Considering that the higher temerature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
-We now return to the hysteretic metamagnetic transitions observed in Figure 3(b).
-Closer inspection into the second metamagnetic transition reveals a complex magnetic landscape which can only be elucidated with high density phase maps.
+Contributions to the magnetic entropy gradually continue up to 100 K, however, consistent with observations in other LnTi3Bi4 materials.[31] By 100 K we recover approximately 90% of the total entropy expected for Tb3+.
+Considering that the higher temperature heat capacity data tends to be noisier (and that the nonmagnetic subtraction is more prone to error at high temperatures), this is remarkably good agreement with expectations.
+We now return to the hysteretic metamagnetic transitions observed in Figure 3(b) and the brief plateau near 2/3 M (Figure 3(b,inset)).
+Closer inspection of the second metamagnetic transition (H 2.8–3.0 T) revealed a much more complex magnetic landscape than the 1.8 K isothermal curve would suggest.
 Figure 4(a,d) present our efforts to map the magnetic landscape of TbTi3Bi4, demonstrating a set of isothermal magnetization curves collected in both the increasing (a) and decreasing (d) field directions.
-One can immediately see that the sharpness of the second transition belies a network of subtle changes in the magnetization.
-Taking the derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e).
-Immediately one can see that the plateau visible in the 1.8 K magnetization data shown before in Figure 3(b) fragments into a multitude of unique states at intermediate temperatures from 2–15 K.
-To better highlight the intricacy of the magnetic landscape near the potential 2/3 magnetization plateau, we collected higher density data over the limited ranges highlighted by the black dashed lines.
-Figure 4(c,f) demonstrate the phase diagram for TbTi3Bi4 in the vicinity of the second metamagnetic transition.
-At least 5 unique phase pockets can be observed.
-Furthermore, while the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field.
-Data was also collected for the zero-field-cooled virgin curves, though they are qualitatively similar to Figure 4(a,b,c) and have been omitted for brevity.
-A series of line cuts (Figure 4(c,right)) through the phase diagram data (white dashed lines) helps to clarify the clear boundaries between the phase pockets.
+As mentioned previously, care must be taken to account for the thermal and magnetic history of TbTi3Bi4 samples.
+Each isothermal curve was performed by warming the sample above 50 K, and then cooling to the desired temperature under zero-field conditions (ZFC).
+As such, the magnetic history is erased for each measurement.
+No ZFC virgin magnetization curves are shown here, but they have been included in the ESI for completeness.
+The sharpness of the second transition belies a network of subtle changes in the magnetization.
+These transitions are quantitatively consistent between samples and between sample batches, suggesting that the phenomenon observed in Figure 3 and Figure 4 are intrinsic to TbTi3Bi4.
+Taking the field-derivative of the magnetization data shown in (a,d) yields the phase diagrams shown in Figure 4(b,e).
+Even at this scale, the plethora of phase pockets near 2/3 M is immediately evident, though they are most pronounced at intermediate temperatures from 2 K–15 K.
+To better highlight the intricacy of the magnetic landscape near the 2/3 M plateau, we collected higher density data (Figure 4(c,f)) over limited temperature [2 K–18 K] and field [2 T–3.5 T] ranges.
+The magnetic history was erased between each isothermal curve as described previously.
+ZFC virgin curves have been omitted, though they have been included in the ESI.
+While the number of pockets appears to be roughly conserved, the qualitative behavior of the diagram depends strongly on whether the data is collected upon increasing or decreasing field.
+A series of line cuts at various temperatures (white dashed lines) have been included alongside both phase diagrams to help visualize the phase boundaries.
 Considering the complexity of the magnetic landscape, neutron diffraction experiments will be invaluable to identify the types and natures of the individual magnetic phases.
-Antiferromagnetic metals are not particularly common, particularly so for materials that can be readily exfoliated.
-While the low temperature magnetism of the rare-earth metals prevents facile applications within fields like spintronics, understanding how the magnetism interacts with the transport properties is key to understanding the connection between the kagome network (which dominates the electronic states near E) and the magnetism.
-Figure 5(a) shows a basic resistivity curve as a function of temperature under zero-field conditions.
-Our experiment passes current parallel to the [010] direction, and orients the field such that (oriented to induce the metamagnetic transitions).
+Considering the complexity of the magnetic landscape, we were curious whether the features observed in Figure 3 and Figure 4 would have a dramatic influence on the magnetotransport.
+Cleavable antiferromagnetic materials have recently been a focus in fields like spintronics,[57, 58] and while the low-temperature magnetism of the rare-earth metals may prevent immediate practical application of these materials, there is still insight to be gleaned from the transport data.
+In particular, probing the connection between complex magnetic phases (via the supporting sublattice) and the transport properties (dominated by the kagome Ti states at E) provides insight into how the sublattices interact.
+Figure 5(a) shows the zero-field, in-plane resistivity of TbTi3Bi4 with (8 mA).
+A simplified schematic of our sample setup is shown in the upper corner of Figure 5(a).
+While the schematic draws the hexagonal faceted crystal to highlight the [010] facet (edge), the actual sample employed in the measurements was a bar cut from the bulk crystal with a length of 0.4 mm and a cross-sectional area of 0.05 mm2.
 The crystal quality is good (RRR40) and the zero-field transport is indicative of a relatively high mobility metal.
 The magnetic transition at 20 K is evident in the resistivity, highlighted by a sharp cusp and subsequent drop.
-This is the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
-We now apply a magnetic field to examine the magnetoresistance (MR) of TbTi3Bi4 in the same experimental setup.
-Figure 5(b) shows the symmetrized transverse magnetoresistance with and at 1.8 K.
-As one would expect from the magnetization results, we observe two sharp transitions in the magnetoresistance, one for each of the metamagnetic features.
-The magnetoresistance retains the qualitative field hysteresis of the magnetization data, both in magnitude and location.
-The first metamagnetic transition marks a sharp increase in the magnetoresistance of approximately 30–40%.
-The second metamagnetic transition marks a sharp decrease in the magnetoresistance to approximately 50%, which one may expect for a total polarization of the Tb moments.
+This is perhaps the expected result for ordering from the paramagnetic regime, sharply reducing spin-scattering.
+We now apply a magnetic field () to examine the in-plane (transverse) magnetoresistance (MR) of TbTi3Bi4.
+Figure 5(b) shows the symmetrized transverse MR with and (8 mA) at 1.8 K.
+As before, we are primarily interested in the magnetotransport under repeated field cycles and have omitted the ZFC virgin curves.
+Instead, these measurements begin in the fully polarized (9 T) state, ramping to 9 T, and finally back to 9 T.
+Symmetrized curves corresponding to increasing field direction (in both positive and negative directions) have been indicated by solid purple lines, whereas the decreasing field direction is represented by dashed purple lines.
+As one would expect from the magnetization results, we observe two sharp transitions in the MR, one for each of the metamagnetic features.
+The first metamagnetic transition marks a sharp increase in the MR of approximately 30–40%.
+The second metamagnetic transition marks a sharp decrease in the MR to approximately 50%, which one may expect for a total polarization of the Tb moments.
 We find the sharp switching between positive and negative MR intriguing in TbTi3Bi4, particularly for a system which is already quite metallic (1Ohm-cm at 1.8 K).
-Given the strong MR response in TbTi3Bi4, we were curious to investigate whether the complex landscape of magnetic transitions seen in the phase diagrams would influence the electronic transport.
-Figure 5(c) provides a series of magnetoresistance curves as a function of temperature from 2 K to 21 K in 1 K steps.
-For simplicity and clarity, we are only showing the data under increasing fields.
-We would like to highlight the behavior near the second transition of the symmetrized data (3 T).
-Immediately before the transition to negative magnetoresistance, we observe that the magnetoresistance cusp becomes a peak – then splits into two peaks – merge back together – and then vanish as the system approaches the Neel temperature.
-These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4(c).
+The strong MR response in TbTi3Bi4 at 1.8 K highlights the sensitivity of the transport to the particular magnetic state.
+Given the complexity of the magnetic phase pockets near the plateau at 2/3 M (see Figure 4(c,f)), we were curious whether the electronic transport would be sensitive to the subtle shifts in the bulk magnetization.
+Figure 5(c) provides a series of MR curves as a function of temperature.
+For simplicity and clarity, we are only showing the data under increasing fields, though full field loops are shown in the ESI.
+The field range corresponding to the 2/3 M plateau (see Figure 3(b)) has been accentuated by the light gray background.
+Immediately before the transition to negative MR, we observe that the MR cusp becomes a peak – then splits into two/three peaks – merge back together – and then vanish as the system approaches the Néel temperature.
+These effects occur at the intermediate temperatures consistent with the appearance of the various phase pockets in Figure 4.
 This further suggests that these are bulk effects corresponding to distinct types of magnetic order.
-The sample geometry prevents us from measuring the Hall response where and .
-The in-plane moments would require thinning the sample perpendicular to the natural layers of the crystal, which typically causes the sample to flake and fall apart.
-However, we can readily orient the sample such that and to explore the Hall effect and magnetoresistance when the metamagnetic transitions are not induced.
+Utilizing the same sample geometry as Figure 5(a–c), we turned to investigate the Hall effect measurement with and .
+It is critical to note that there are no metamagnetic transitions with , so both the MR and Hall effect are probing a much simpler magnetic landscape.
+Furthermore, the effect of hysteresis (and the sensitivity to field and thermal history) vanishes.
 Figure 5(d) shows the antisymmetrized Hall resistivity for TbTi3Bi4 over two temperature regimes.
-The top figure shows the Hall response from 2 K to 18 K, and the bottom figure highlights temperatures above the magnetic transition temperature (20 K to 30 K).
+The top figure shows the Hall response within the antiferromagnetic state (2 K – 18 K), and the bottom figure highlights temperatures within the paramagnetic phase (20 K – 30 K).
+As with the MR, full field loops were collected at each temperature, starting at the fully-polarized 7 T state, ramping to 7 T, and finally back to 7 T.
 No subtractions or corrections have been performed besides antisymmetrization.
-All data sets exhibit a strongly nonlinear Hall response over all fields.
-A signal that looks deceptively like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature.
+All data sets shown here exhibit strongly non-linear behavior.
+A signal that looks curiously like an anomalous Hall signal is evident at 2 K, though the signal persists in shape and magnitude above the magnetic transition temperature.
 Furthermore, at high temperatures we recover an additional inflection in the data.
 The signal is also relatively small, leading us to suspect that we may be seeing the effect of multiband transport.
-We can look to some prior work on NdTi3Bi4 for some guidance – since the electronic structure of these materials doesn’t appreciably change with the choice of rare-earth.
-Experimental ARPES data has observed a Dirac-like feature near the Fermi level[38, 37, 35, 36] and has further observed gapping of the Dirac-like bands as NdTi3Bi4 transitions into the magnetic ground state[35].
-The sample exhibits strong positive magnetoresistance (200%) when and , in stark contrast to behavior when and .
+Finally, Figure 5(e) presents the transverse MR when and .
+In the absence of the metamagnetism, no hysteresis in the MR is observed.
+Further, the system exhibits only strong positive MR (200%) over all fields and temperatures.
+Nearly simultaneously with this work, a manuscript focusing on the magnetotransport and ARPES in TbTi3Bi4 were posted.[54] This work complements the results shown here by focusing on the Hall effect when (inducing the metamagnetic transitions).
+The authors note an record-high anomalous Hall conductivity, nearly an order of magnitude larger than comparable systems.
 Overall, our work into TbTi3Bi4 has revealed a unique antiferromagnetic, cleavable metal with an exceedingly complex landscape of magnetic phases.
-The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative magnetoresistance near the 2/3 magnetization plateau, and an unusual coupling between the magnetic phase pockets and the transport.
+The magnetic and transport properties are rich and diverse, with multiple unique magnetic phases to investigate, strong switching from positive to negative MR near the 2/3 M plateau, and a complex coupling between the magnetic phase pockets and the transport.
 We now turn to briefly investigate the magnetic properties of the Ln2-xTi6+xBi9 shurikagome metals and the influence of disorder on the Ln–Ln chains.
-Shurikagome Ln | Occ | Occ | Tot Ln | Tot Ln | M () | M () | M () | T (K) | () | (K) |
-Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.7 | -24 |
-Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | -14 |
-Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | -0.2 |
-Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.6 | -0.9 |
-Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 7.3 | 14 |
-Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | 0 | 0 | 0.1 | N/A | N/A | N/A |
-III.3 Properties of Ln2-xTi6+xBi9
-Diffraction studies on the Ln2-xTi6+xBi9 compounds revealed several interesting observations.
-We first attempted to solve the structure using fully occupied Ln sites (forming fully occupied chains).
-While R values were actually quite favorable, there existed large amounts of excess electron density around the Ln04 and Ln05 sites.
+III.3 Properties of Ln2-xTi6+xBi9 (Ln: Tb–Lu) family
+When we first identified the Ln2-xTi6+xBi9 shurikagome metals, we were surprised by the substitution of Ln into the Ti8Bi9 structure.
+Unlike integration of the rare-earth elements into frameworks like CoSn, there are no clear interstitial voids in Ti8Bi9.
+Further, the miscibility of most rare-earth elements into Ti is low, and it is initially unintuitive that Ln is the predominant defect.
+The incomplete substitution of Ln is also perplexing.
+While single crystal X-ray diffraction (SCXRD) results can be solved using fully occupied Ln sites (forming fully occupied chains) with relatively favorable statistics, large amounts of excess electron density accumulate on the Ln4 and Ln5 sites.
 The residual Q-peaks can be eliminated by allowing the occupancy of the Ln sites to vary, which results in partially occupied Ln chains.
-However, compositions extracted from single-crystal X-ray diffraction (SCXRD) using the assumption of pure (although partially occupied) Ln sites are in strong disagreement with electron dispersive spectroscopy (EDS) results.
-Bluntly, there is not enough Ti in the unit cell derived from SCXRD.
-This can be remedied by allowing antisite disorder on the Ln sites, allowing the composition of the Ln to vary, but restraining it such that the site remains fully occupied (e.g.
-Ln04 + Ti04 = 1).
+This is also consistent with energy dispersive spectroscopy (EDS) results, which support a Ln content from 6–9% (ideal Ln2Ti6Bi9 is 12% Ln).
+For simplicity, we have restrained the chain sites to remain fully occupied (e.g.
+Occ + Occ = 1).
 Even this appears to slightly underestimate the amount of Ti, and EDS results suggest that some Ti may be substituting elsewhere as well.
-This is perhaps expected, as the published Ti–Bi phase diagram indicates a full solid solution over the range of 50:50 Ti:Bi to 35:65 Ti:Bi.[41]
-However, the SCXRD data did not independently suggest any additional site substitutions, and the constraints mentioned above yielded quite favorable statistics.
+The published Ti–Bi phase diagram supports Ti substitution, as Ti8Bi9 supports a full solid solution from 50:50 Ti:Bi to 35:65 Ti:Bi.[51] Regardless, the SCXRD data did not improve substantially with more complex substitution models.
+SCXRD tables, CIF files, and EDS results for the entire series can be found in the ESI.
+The substitution of the Ln atoms creates a clear separation between the Ln4/Ln5 sites and the Ti-based shurikagome network.
+As suggested by our designation “shurikagome,” the coordination of the Ti sublattice in the Ln2-xTi6+xBi9 compounds contains pieces of the corner-sharing triangular motif observed in prototypical kagome samples.
+The parent structure Ti8Bi9 has been known for some time,[50] though the electronic structure has only been calculated through high-throughput compendiums like Materials Project.[59] In an effort to understand the underlying electronic structure of the Ln2-xTi6+xBi9 compounds, we have performed first-principle density functional theory calculations on the ideal Ln2Ti6Bi9 structure.
+To avoid the additional complication of magnetism, we examine the electronic structure of Lu2Ti6Bi9 as a proxy for the material family.
+Figure 6 provides the calculated electronic structure for Lu2Ti6Bi9 alongside a schematic diagram of the P4/nmm Brilluoin zone.[60] There are several interesting features apparent within the data.
+We first note a Fermi-level band edge and associated “flat–band,” in two distinct directions, centered at ().
+The extent of this flat band is relatively limited when compared to the prototypical kagome band structure, though the proximity to the Fermi level is interesting.
+Similar to the case of the LnTi3Bi4 compounds, the states near the Fermi level are almost entirely comprised of Ti-based orbitals arising from shurikagome motif.
+Beyond the flat-bands, the Lu2Ti6Bi9 band structure contains many dispersive Dirac-like features and several avoided crossings near E (at and near the -point, 10 meV above E between and , 50 meV above E between and ).
+Flat bands and/or van Hove singularities near E are commonly cited as the cause for a wide array of unusual behavior, presumably due to the localized nature of electrons in flat-bands and associated correlative effects.
+The numerous band edges at or very near (25 meV) the Fermi level in Lu2Ti6Bi9 are suggestive of numerous Fermi surface Lifshitz transitions, which supports the off-stoichiometric behavior of these compounds.
+This additionally indicates a complex, highly sample-dependent Fermi-surface structure.
+Further, while the spectral weight associated with the Fermi-level “flat-bands” in Lu2Ti6Bi9 is relatively small, it is still suggestive of potential novel behavior potentially observable via ARPES.
+Out of curiosity, we performed preliminary ARPES experiments on an in-house laser ARPES system on crystals of Tb2Ti6Bi9 (which has the highest rare-earth occupancy) at approximately 10 K.
+We were able to observe a high density of states near the Fermi level, consistent with Figure 6 and the associated “flat-bands.” However, the intrinsic disorder in the Ln2-xTi6+xBi9 compounds is currently prohibitive to performing in-depth ARPES investigations, particularly considering the relatively complex electronic structure.
+Whether the flat bands can be directly ascribed to the Ti-based “shurikagome” sublattice (in analogy to tight binding calculations in the prototypical kagome metals) is a subject of current investigation.
+We believe that the applicability of both the DFT and ARPES models will require parallel efforts to synthesize compounds with reduced disorder.
+The existence of the halfnium-based Hf8Bi9[50] and some similar vanadium-based V8Sb9[61] compounds may offer some flexibility in the choice of framework cations that may offer routes to achieve fully-occupied Ln2M6X9 samples.
+Indeed, synthesizing fully-occupied Ti8Bi9-based shurikagome compounds may be challenging, as the extent of disorder in samples of Ln2-xTi6+xBi9 appear to vary quite strongly with the rare-earth species.
 Figure 7(a) highlights the cell volume for the Ln2-xTi6+xBi9 shurikagome metals as a function of the 8-coordinate Shannon radius.
@@
 These results are briefly summarized in Table 1.
-While CIF files are provided in the ESI, here we focus on the influence of the Ln atom and the disorder in crystals of Ln2-xTi6+xBi9.
-The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln04 and Ln05.
-As mentioned before, the Ti occupancy on those sites is constrained such that they add to 1.
+The columns labeled Occ and Occ indicate the refined occupancy of the chain sites Ln4 and Ln5.
 The sum of the SCXRD-determined occupancies is provided in the column labeled “Tot Ln.” This can be directly compared with the composition extracted from EDS measurements in the next column.
-SCXRD appears to consistently calculate higher concentrations of Ln as compared to EDS measurements, over approximating by about 15%.
-Some interesting trends appear in the overall data.
-Both the Ln04 and Ln05 site occupancies decrease as we move to smaller Ln atoms.
-This effect is substantially more pronounced on the Ln05 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9.
+Some interesting trends appear in the agglomerate data.
+First, SCXRD appears to consistently calculate higher concentrations of Ln when compared to EDS.
+However, regardless of technique, both the Ln4 and Ln5 site occupancies decrease as we move to smaller Ln atoms.
+This effect is substantially more pronounced on the Ln5 (smaller) rare-earth site, where the occupancy is nearly halved by the time we reach Lu2-xTi6+xBi9.
 We performed some exploratory syntheses around Tb2-xTi6+xBi9 in an attempt to fully saturate the Ln site, though only small changes in the rare-earth composition were noted.
-Considering the high disorder in crystals of Ln2-xTi6+xBi9, particularly within the Ln sublattices, it is not clear what magnetic response to expect.
-Naively, the most likely scenario was freezing into a spin-glass, or a complete loss of a well-defined magnetic ground state in the non-Kramers ions.
-We now turn to a brief overview of the magnetic properties of the Ln2-xTi6+xBi9 compounds.
-The compounds are tetragonal (P4/nmm) yielding two directions of interest, (parallel to chain) and (perpendicular to chain).
-We did not consider rotations with in the ab plane for this work.
+Coincidentally, we have not observed the formation of Ln2-xTi6+xBi9 with large rare-earths (La–Gd).
+We speculate that the larger rare-earths may destabilize the structure.
+However, under this assumption, we find it unusual that smaller rare-earth elements have decreasing occupancies, particularly on the smaller site.
+Remarkably, despite the high degree of chemical disorder in crystals of Ln2-xTi6+xBi9, magnetization measurements exhibit properties more indicative of antiferromagnetic order.
+Due to the tetragonal (P4/nmm) structure and the quasi-1D Ln–Ln chain, there are two directions of interest for magnetization measurements, (parallel to chain) and (perpendicular to chain).
+Samples containing Tb, Dy, and Ho exhibit a cusp in the magnetization when , whereas Er and Tm have features when instead.
 All compounds exhibit weak anisotropy with the exception of Tm.
 When , Tm2-xTi6+xBi9 exhibits an order of magnitude stronger response.
-For Tb, Dy, and Ho, data collected with exhibits a cusp in the magnetization, strangely resembling an antiferromagnetic transition.
-For Er and Tm, the data with exhibits a broad peak instead.
-All other combinations of elements and directions (e.g.
-Tb with are featureless, although the magnitude of the magnetization remains comparable to the other direction).
-These results are summarized in Figure 6(a) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction.
-Please note that the magnetization results are shown in arbitrary units for ease of comparison.
-Full data sets are provided in the ESI.
-Note that the data presented in Figure 6(a) includes both zero-field-cooled (ZFC) and field-cooled (FC) data.
+These results are summarized in Figure 7(b) which provides an overview of the magnetization for the Ln2-xTi6+xBi9 shurikagome metals along their “favored” direction.
+Note that Figure 7(b) is shown in arbitrary units for ease of comparison, though full data sets are provided in the ESI.
+Shurikagome Ln | Occ | Occ | Tot Ln | Tot Ln | M () | M () | M () | T (K) | () | (K) |
+Tb2-xTi6+xBi9 | 0.89 | 0.76 | 1.65 | 1.4 | 7.4 | 6.3 | 7.0 | 11.5 | 9.8 | 21 |
+Dy2-xTi6+xBi9 | 0.84 | 0.62 | 1.46 | 1.3 | 6.8 | 5.9 | 6.8 | 7.8 | 9.6 | 14 |
+Ho2-xTi6+xBi9 | 0.83 | 0.56 | 1.39 | 1.2 | 6.9 | 6.0 | 6.9 | 3.2 | 9.9 | 0.8 |
+Er2-xTi6+xBi9 | 0.81 | 0.54 | 1.35 | 1.2 | 6.1 | 5.3 | 6.2 | 2.4 | 10.5 | 1.0 |
+Tm2-xTi6+xBi9 | 0.81 | 0.56 | 1.37 | 1.2 | 4.8 | 4.3 | 5.3 | 1.6 | 8.6 | 1.1 |
+Lu2-xTi6+xBi9 | 0.75 | 0.46 | 1.21 | 1.0 | N/A | N/A | N/A | N/A | N/A | N/A |
+The data presented in Figure 7(b) includes both ZFC and FC data for Tb–Er samples.
 The lack of ZFC/FC splitting is surprising, as we may reasonably expect the strong disorder to result in a glassy magnetic state.
 Furthermore, AC susceptibility measurements on Tb2-xTi6+xBi9 indicate no frequency dependence.
-However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits a strong frequency response.
-It is striking that the features follow the temperature scale expected of the latter half of the rare-earth row – Tb possesses the highest ordering temperature, and Tm possesses the lowest.
-Figure 6(b) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K.
-Again we have opted to show only one direction (the same direction as in Figure 6(a)) for simplicity.
+However, this is not universal throughout the series, as Tm2-xTi6+xBi9 exhibits some frequency dependence.
+Notably, however, Tm also possesses one of the lowest Ln occupancies (Tm1.4Ti2.6Bi9) across the series.
+The magnetization anomalies follow the temperature scale (de Gennes scaling) expected of the latter half of the rare-earth row, with Tb possessing the highest ordering temperature, and Tm possessing the lowest.
+The results are generally reminiscent of antiferromagnetic-like order, particularly considering the lack of ZFC/FC splitting.
+As an additional diagnostic tool, we performed zero-field heat capacity measurements (see ESI) on both Tb2-xTi6+xBi9 and Er2-xTi6+xBi9 crystals from 300 K to 1.8 K.
+Both samples show clear, albeit small, heat capacity anomalies.
+These results are in good agreement with magnetization results and further support some degree of non-glassy magnetic order in the series.
+Figure 7(c) demonstrates the isothermal magnetization for all magnetic Ln2-xTi6+xBi9 compounds at 1.8 K.
+For simplicity we have opted to show the same field directions as in Figure 7(b), though both the and directions are included in the ESI.
 All compounds near saturation by 7 T, though Er2-xTi6+xBi9 has not saturated, owing to an unusual inflection in the data near 4 T.
 Regardless, we can extract the approximate saturation magnetization for all compounds, which has been summarized in Table 1 (column M).
-Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2-xTi6+xBi9.
-This will underestimate the true saturation magnetization, as we will be normalizing to 2 rare-earth atoms per formula unit instead of (2-x).
-However, we can compare these results with the expected saturation magnetization from XRD and EDS results (M and M).
-Comparing directly, we find that the experimental results actually agree quite well with the SCXRD results.
+Our quantitative analysis assumes the molar susceptibility can be expressed in terms of the ideal formula unit Ln2Ti6Bi9.
+This underestimates the true saturation magnetization, though it can be compared directly with the “expected saturation magnetization from XRD and EDS results (M and M).
 Some other parameters of interest (e.g.
 and ) have also been added to Table 1.
-Due to the disorder and variable number of Ln sites, we consider these values as first approximations.
-Despite the high levels of disorder, some samples of Ln2-xTi6+xBi9 (e.g.
-Tb) appear to show magnetization consistent with long-range order.
-This is highly unexpected, and considering the lack of ZFC/FC splitting and no clear frequency dependence, more detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed.
+Of course, due to the disorder and variable number of Ln sites, we consider these values as first approximations.
+Remarkably, despite the high levels of disorder, the majority of Ln2-xTi6+xBi9 appear to show magnetic properties consistent with non-glassy magnetic order.
+The lack of ZFC/FC splitting, the lack of frequency dependence in most samples, and the presence of well-defined heat capacity anomalies suggest some degree of conventional magnetic ordering.
+The disorder assuredly plays a non-trivial role in the manifestation of this order, and we cannot completely exclude the possibility of some glassiness.
+More detailed investigations into the magnetism and chemical disorder of the Ln2-xTi6+xBi9 family are clearly needed.
+Future work should focus on controlling chemical disorder in the shurikagome metals, particularly with an emphasis on realizing an idealized Ln2M6X9 analog.
 IV Conclusions
-We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, a kagome metal belonging to the AM3X4 family of materials.
-We described the rational for the formation of TbTi3Bi4, including the increased competition with another class of materials, the Ln2-xTi6+xBi9 shurikagome metals for small rare-earth species (Ln: Tb–Lu).
-Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases surrounding multiple metamagnetic transitions.
+We began this work by describing a modified synthesis procedure that is capable of subverting phase competition to realize TbTi3Bi4, the smallest rare-earth containing member of the LnTi3Bi4 kagome metals.
+We described the rationale for the formation of TbTi3Bi4, including the increased competition with the shurikagome metals Ln2-xTi6+xBi9 (Ln: Tb–Lu).
+Our work into both classes of materials have demonstrated that TbTi3Bi4 is an exceedingly interesting candidate material – an exfoliable, antiferromagnetic kagome metal with a complex landscape of magnetic phases, metamagnetic transitions, and 1/3 and 2/3 magnetization plateaus.
 Transport measurements highlight the impact the rare-earth magnetism has on the kagome-dominated transport.
-Upon entering the magnetization plateau, the system sharply switches between positive (30-40%) and negative (-55%) magnetoresistance in a system that is already quite metallic (1 Ohm-cm).
+Upon entering the magnetization plateau, the system sharply switches between positive (+40%) and negative (50%) magnetoresistance in a system that is already quite metallic (1 Ohm cm).
 We have also briefly examined the properties of the Ln2-xTi6+xBi9 shurikagome metals.
-Despite considerable disorder on the rare-earth sublattices, some systems exhibit magnetization responses more indicative of long-range order.
-Together, these results not only provide insight into a new kagome lattice and an unusual family of disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
+Despite considerable disorder on the rare-earth sublattices, the systems exhibit magnetization responses reminiscent of non-glassy antiferromagnetic order.
+Potential routes towards fully-occupied Ln2M6X9 analogs may provide a cleavable system with an interesting electronic structure.
+Together, our results not only provide insight into a new kagome lattice and an unusual family of chemically-disordered magnets, but also illuminate a synthetic strategy to help realize new members of the AM3X4 family.
 V Acknowledgements
-The work of B.R.O (synthesis, data collection, writing), H.Z.
-(analysis), K.G.
-(He3 data collection), A.F.M.
-(He3 data collection), and M.A.M.
-(editing), was supported by the U.S.
+This work was supported by the U.S.
 Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division.
-We thank the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source for use of their Photonics Laue Camera System.
+We thank Jong Keum and the X-ray laboratory of the Oak Ridge National Laboratory Spallation Neutron Source (SNS) for use of their Photonic Science Laue camera.
+Research directed by G.D.S.
+is additionally sponsored by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy.
+This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S.
+Department of Energy under Contract No.
+DE-AC05-00OR22725.
+Portions of this work were supported by the U.S.
+Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (ARPES, Q.L., R.G.M.)
+VI Supplementary Info.
+Supplementary crystallography tables and corresponding CIF files; Additional magnetic phase diagram utilizing zero-field-cooled virgin curves; Preliminary ARPES data on TbTi3Bi4; Additional magnetoresistance data with full temperature-field data set and full field loops; Table of shurikagome compositions measured by EDS; Preliminary ARPES data on Tb2-xTi6+xBi9; Magnetization versus temperature data for Ln2-xTi6+xBi9 to scale; Isothermal magnetization data for Ln2-xTi6+xBi9; Curie-Weiss fits for Ln2-xTi6+xBi9; Select heat capacity results for Tb2-xTi6+xBi9 and Tm2-xTi6+xBi9